OpenStudy (anonymous):

Evaluate the integral of cos(x) / 3-sin(x) I'm stuck...

OpenStudy (anonymous):

i'll type my working out

hartnn (hartnn):

did you try the u substitution as \(\large u= 3-\sin x\)

OpenStudy (anonymous):

yep

hartnn (hartnn):

du =... ?

OpenStudy (anonymous):

du/d = -cos(x)

OpenStudy (anonymous):

dx = du / -cos(x)

hartnn (hartnn):

so, -du = cos x dx right ?

OpenStudy (anonymous):

don't you solve for dx though?

hartnn (hartnn):

you can...not necessary...

OpenStudy (anonymous):

u = 3-sin(x) \[\frac{ du }{ dx }=-\cos(x)\] \[dx = \frac{ du }{ -\cos(x) }\]

OpenStudy (anonymous):

well that's the way amistre showed me and also I copied some notes in a lecture today which follows that way as well

hartnn (hartnn):

yes, do that then, and tell me what yur new integral is

OpenStudy (anonymous):

\[\int\limits_{}^{}-\frac{ 1 }{ u }du\]

hartnn (hartnn):

correct, know the standard formula for that ?

OpenStudy (anonymous):

that's after i substitute the u and cancel out the cos(x)

OpenStudy (anonymous):

nope, I'm stuck at that part :(

OpenStudy (anonymous):

wait is it -log(u) + C?

hartnn (hartnn):

\(\large \int \dfrac{f'(x)}{f(x)} dx= \log f(x)+c \\ \large \int \dfrac{1}{x} dx= \log x+c\) yes!

OpenStudy (anonymous):

so the final answer should be -log(3-sin(x))+C i believe?

hartnn (hartnn):

correct! :)

OpenStudy (anonymous):

k thanks!

hartnn (hartnn):

welcome ^_^